OpenStudy (anonymous):

A system of equations is shown below: 6x − 5y = 5 3x + 5y = 4 The x-coordinate of the solution to this system of equations is _____. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (dbzfan836):

@_adore.ki , do you understand what to do inside of this problem?

OpenStudy (dbzfan836):

If not, we're gonna take a bit of time to explain what has to be done to solve.

OpenStudy (anonymous):

no i don't understand

OpenStudy (dbzfan836):

Alright then. Do you want the long but easy method or annoying and also easy method? XD

OpenStudy (anonymous):

lol the long but easy method

OpenStudy (dbzfan836):

alright, this method just needs a pen and paper.

OpenStudy (dbzfan836):

You see, x is the input, and y is the output

OpenStudy (dbzfan836):

All inputs and going to be the same in these situations, but the outputs different.

OpenStudy (dbzfan836):

6x − 5y = 5 3x + 5y = 4 We need to find an x that would work.

OpenStudy (dbzfan836):

So let's say x is 1. 6(1) − 5y = 5 3(1) + 5y = 4

OpenStudy (dbzfan836):

What would y be for the equations?

OpenStudy (anonymous):

I really don't know :(

OpenStudy (dbzfan836):

hmm. Let's do this one together then. 6(1) − 5y = 5 6 - 5(.2) = 5 6 - 1 = 5 3(1) + 5y = 4 3 + 5(.2) = 4 3 + 1 = 4 As you can see, we have already arrived at the answer. However, things are not always so simple

OpenStudy (dbzfan836):

The answer is 1.

OpenStudy (anonymous):

for this whole question the answer is 1 ?

OpenStudy (dbzfan836):

the solution for a system for equations is the one point where the two lines meet. both of the equations are linear, meaning they will always be in a straight line on a graph. if the x and y are same, then they have made their one and only intersection with the other equation. Thus, the solution has been found.

OpenStudy (anonymous):

So the answer was 1 ?

OpenStudy (dbzfan836):

Yes.

OpenStudy (anonymous):

Thank you